OpenStudy (anonymous):

A potato with temperature 70 degrees F is placed into a oven with an internal temperature 400 degree after 30 minutes, the temperature of the potato has increased to 205 degrees F. Give the heating constant k (rounded to 6 decimal places), and find an expression for the temperature potato as a function of time.

OpenStudy (nottim):

so useful... whoa...was gunna give the anseer there for some reason...

OpenStudy (anonymous):

huh?

OpenStudy (nottim):

well, what's our initial constant? What would the entire thing equal?

OpenStudy (anonymous):

A potato with temperature 70 degrees F is placed into a oven with an internal temperature 400 degree. After 30 minutes, the temperature of the potato has increased to 205 degrees F.

OpenStudy (nottim):

uh

OpenStudy (anonymous):

You need the Newton's formula for this one.\[T(t)=T_1+(T_0-T_1)e ^{-kt}\]T0 is the starting temperature of the object, T1 is the temperature of the environment.

OpenStudy (anonymous):

As t gets larger, the second term fades away, which leads to the temperature of the object T(t) being T1 (the temperature of the surroundings.

OpenStudy (anonymous):

So plug the data from the problem into the formula; you will have only one unknown value (k) that can be solved for.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Solve this....\[205=400+(70-400)e^{-30k}\]

OpenStudy (anonymous):

what is the temperature of the potato after 80 minutes?

OpenStudy (anonymous):

After what amount of time will the temperature of the potato reach 350 degrees F?

OpenStudy (anonymous):

You can find that after you have solved for k.

OpenStudy (anonymous):

The last unknown is k; once you have that, you just plug the time or temperature into the formula, and there will be only one unknown to solve for.

OpenStudy (anonymous):

\[205=400+(70−400)e^{−30k}\implies \frac{-195}{-330}=e^{−30k}\]\[\implies \frac{\ln(\frac{195}{330})}{-30}=k\]

OpenStudy (anonymous):

After you have k (use your calculator to get an approximation), you can now solve for unknown time or temperature easily.

OpenStudy (anonymous):

so 80 replaces 205 in the equation to solve for the temperature after 80 minute?

OpenStudy (anonymous):

No, you put 80 in for t, the time in the exponent. T(t) is the temperature of the object as a function of t (time). Previously, the time was 30 in the problem above. See it?